OpenStudy (anonymous):

C=2 pi ''aka 3.14'' r

OpenStudy (anonymous):

Solve for r

OpenStudy (anonymous):

c = 2(3.14)r r = c/2(3.14) r = c/6.28 Answer: r = c/6.28